OpenStudy (hang254):

A circle is represented by the equation shown below. (x - 5)^2 + (y + 9)^2 = 144 Which statement is true? The circle is centered at (-5, -9) and has a diameter of 12. The circle is centered at (5, -9) and has a radius of 12. The circle is centered at (5, 9) and has a diameter of 12. The circle is centered at (5, 9) and has a radius of 12.

OpenStudy (hang254):

@campbell_st

OpenStudy (campbell_st):

ok... the general form for a circle is \[(x - h)^2 + (y - k)^2 = r^2\] and centre at (h, k) and radius r... you you use this to find the centre and radius of your circle..?

OpenStudy (hang254):

so i substitute the x and y values in to the equation?

OpenStudy (campbell_st):

nope... just what do you think the value of h is... when you compare the equations..?

OpenStudy (hang254):

h=x value?

OpenStudy (campbell_st):

no... have a look (x - h)^2 (x - 5)^2 what is the value of h in your equation?

OpenStudy (campbell_st):

compare (x - h)^2 + (y - k)^2 = r^2 general form (x - 5)^2 + (y + 9)^2 = 144 Your question any ideas for h, k or r